OpenStudy (anonymous):

Find the general solution of the equation (y^2 cosx-3x^2y-2x)dx+(2ysinx-x^3+lny)dy=0

OpenStudy (anonymous):

My 1st try is 2 show this diffenrential equation is an exact equation. So,\[M =y ^{2}\cos x -3x ^{2}y -2x\] \[\frac{ dM }{ dy }= 2ycox-3x ^{2}\] \[N =2ysinx -x ^{3}+ \ln y\] \[\frac{ dN}{ dx }=2y \cos x -3x ^{2}\] Therefore the differential equation is an exact eqution.

OpenStudy (anonymous):

Next,I integrate M and N respectively where \[U{1}= y ^{2}\sin x -x ^{3}y -2xy +\phi(y)\] \[U{2}=y ^{2}\sin x -x ^{3}y +\frac{ 1 }{ y}+\phi(x)\] Then i compare u1 and u2 in order to get the \[\phi(y) and \phi(x)\]

OpenStudy (anonymous):

So it become \[-2xy +\phi(y)=\frac{ 1 }{ y}+\phi(x)\] Then, what is my x and y?

OpenStudy (shubhamsrg):

why dont you simply just, after your work,do this : dM/dN * dx/dy = 1 or dy/dx = dM/dN your original eqn was Ndy + Mdx =0 => dy/dx = -M/N = dM/dN simpler now ?

OpenStudy (anonymous):

Yeah,its bcome simpler.Bt what ll b the ans if i wanna use the method i did above?

OpenStudy (callisto):

\[(y^2 cosx-3x^2y-2x)dx+(2ysinx-x^3+lny)dy=0\] \[M=y^2 cosx-3x^2y-2x\]\[N=2ysinx-x^3+lny\] \[\frac{\partial M}{\partial y}=2ycosx-3x^2\]\[\frac{\partial N}{\partial x}=2ycosx-3x^2=\frac{\partial M}{\partial y}\]So, the DE is exact. f(x,y) = c is a solution \[\frac{\partial f}{\partial x}=M=y^2 cosx-3x^2y-2x\]\[f(x,y) = \int (y^2 cosx-3x^2y-2x )dx=y^2sinx-x^3y-x^2+\phi (y)\]\[\frac{\partial f}{\partial y}=\frac{\partial }{\partial y}(y^2sinx-x^3y-x^2+\phi (y))=N\]\[2ysinx-x^3+\phi'(y)=2ysinx-x^3+lny\]\[\phi'(y)=lny\]So, you can get \(\phi '(y)\) from here.

OpenStudy (callisto):

Sorry! Last sentence: So, you can get ϕ(y) from here.

OpenStudy (callisto):

Then, \[f(x,y) = y^2sinx-x^3y-x^2+\int lny dy=C\] (Do the integration..)

OpenStudy (anonymous):

Hmm,i see my mistake.Thx ur guys' help :)